4) A 20 year, 4%, $1000 face value bond paying annual coupons is callable at par at the end of any year from 15 on.

a) How much should an investor pay if they want to earn at least 6% until the bond is called (or matures).

b) How much should an investor pay if they want to earn at least 3% until the bond is called (or matures).
Answer #1

1.
=MIN((4%*1000/6%*(1-1/1.06^15)+1000/1.06^15),(4%*1000/6%*(1-1/1.06^20)+1000/1.06^20))=770.601575628695

2.
=MIN((4%*1000/3%*(1-1/1.03^15)+1000/1.03^15),(4%*1000/3%*(1-1/1.03^20)+1000/1.03^20))=1119.37935086776